Circular No. 7164

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990510 J. Kaluzny, Warsaw University Observatory (WUO); P. M. Garnavich

and K. Z. Stanek, Harvard-Smithsonian Center for Astrophysics; W. Pych, WUO; and I. Thompson, Carnegie Observatories, write: “We observed the optical transient found in the error box of GRB 990510 (IAUC 7160) by Vreeswijk et al. (GCN 310) with the 1.0-m Swope telescope at the Las Campanas Observatory between May 11.0 and 11.4 UT. We measure the following precise position for the transient: α = 13^{h}38^{m}07^{s}.64, δ = −80^{o}29′48′′.8 (equinox 2000.0; uncertainty ± 0′′.3), based on the Guide Star Catalog. We find that the object declined steadily in the R band over the observing period and, combined with data from Galama et al. (GCN 313), the light curve is consistent with a power law in time with an exponent of −0.89. There is a suggestion that the slope may be becoming steeper. A comparison star of mag R = 16.5 (position end figures 00^{s}.82, 11′′.7) in the USNO A2.0 catalogue gives R = 18.9 for the optical transient on May 11.0. We find the R magnitudes in this region of the USNO catalogue to be of dubious quality, especially fainter than 16.5. Adding the R-band exposures taken in ∼ 1′′ seeing reveals an object 1′′.8 south of the optical transient and ∼ 3 mag fainter; this may be the host galaxy.”

G. Pietrzynski and A. Udalski, Warsaw University Observatory, report on behalf of the OGLE microlensing survey team: “We monitored the optical counterpart, starting on May 10.992 UT, with the 1.3-m Warsaw telescope at the Las Campanas Observatory. In total, 17 V -band and 21 I-band frames were collected, covering ∼ 10.25 hr (exposure times 900 and 600 s for the V and I-band frames, respectively). The object is apparently variable — its brightness faded by ∼ 0.6 mag in the course of the night, from 19.45 to 20.05 in V and from 18.50 to 19.15 in I. Photometry of the optical counterpart of GRB 990510 and VI light curves are available from

the OGLE Internet archive at http://www.astrouw.edu.pl/~ftp/ogle.”

COMET C/1998 M5 (LINEAR) Visual m_{1} estimates: Apr. 2.76 UT, 9.8 (A. Baransky, Kiev, Ukraine,

0.20-m reflector); 11.82, 9.5 (M. Reszelski, Szamotuly, Poland, 20×60 binoculars); 18.83, 10.6 (M. Meyer, Frauenstein, Germany, 0.25-m reflector); 21.06, 9.6 (K. Hornoch, Lelekovice, Czech Republic, 10×80 binoculars); May 3.90, 11.1 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 9.85, 11.0 (M. Lehký, Hradec Králové, Czech Republic, 25×100 binoculars); 11.10, 11.0 (J. Bortle, Stormville, NY, 0.41-m reflector).

1999 May 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green